OpenStudy (anonymous):

which of the following is the radical expression of 3 sqrt 4n

OpenStudy (saifoo.khan):

\[3\sqrt{4n}\]

OpenStudy (anonymous):

i know lol thats what i wrote

OpenStudy (saifoo.khan):

haha, wht to do now?

OpenStudy (anonymous):

gimme a sec

OpenStudy (saifoo.khan):

\begin{array}l\color{red}{\text{G}}\color{orange}{\text{i}}\color{#9c9a2e}{\text{v}}\color{green}{\text{e}}\color{blue}{\text{n}}\color{purple}{\text{}}\end{array}

OpenStudy (anonymous):

1. (4n)^3 2.3n^4 3.(4n)1/3 4. 4n1/3

OpenStudy (saifoo.khan):

\[\huge \sqrt[3]{4n}\] is it like this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Oh, then the answer is the 3rd one.